# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v3

@@ Abstract
Recent advances in deep learning-based joint source–channel coding (deepJSCC) have substantially improved communication performance, but their high computational cost hinders practical deployment. Moreover, certain applications require the ability to dynamically adapt computational complexity. To address these issues, we propose a Feature Importance-Aware deepJSCC (FAJSCC) model for image transmission that is both computationally efficient and adjustable. FAJSCC employs axis-dimension specialized computation, which performs efficient operations individually for each spatial and channel axis, significantly reducing computational cost while representing features effectively. It further incorporates selective deformable self-attention, which applies self-attention only to selected and adaptively adjusted features, leveraging the importance and relations of input features to efficiently capture complex feature correlations. Another key feature of FAJSCC is that the number of selected important areas can be controlled separately by the encoder and the decoder, depending on the available computational budget. It makes FAJSCC the first deepJSCC architecture to allow independent adjustment of encoder and decoder complexity within a single trained model. Experimental results show that FAJSCC achieves superior image transmission performance under various channel conditions while requiring less computational complexity than recent state-of-the-art models. Furthermore, experiments independently varying the encoder and decoder’s computational resources reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. The code is publicly available at github.com/hansung-choi/FAJSCCv2.
@@ I Introduction
Although conventional separation-based digital communication systems have undergone rapid improvements in both source coding (e.g., JPEG, JPEG2000, PNG, BPG, and AVIF) and channel coding (e.g., Turbo, LDPC, and polar codes) over several decades, further manual improvements to these rule-based methods have become increasingly difficult in satisfying recent communication requirements. In other side, it is known that the joint source-channel coding (JSCC) design strictly outperforms the separation-based design in several scenarios, such as finite block length [1] and sending correlated data over multi-user settings [2]. Early analytic research on the JSCC was limited to simple data distributions, like Gaussian cases, due to the tractability, and hence could not deal with real data, such as images [3, 4].
@@
Beyond such compression-based strategies, another line of work focuses on designing lightweight models [29, 30, 31, 32, 33]. For example, reducing the intermediate deepJSCC feature size before passing through computationally intensive blocks [33] or performing computations only on a fixed subset of feature channels [29, 31] can lower the computational load. In addition, approaches that remove the least important features have been proposed [30, 34]. While these methods alleviate computational burden without requiring extra training, reducing or deleting features and restricting computations to limited channel regions degrade image transmission performance. Therefore, there is an urgent need for deepJSCC designs that can improve transmission performance while reducing computational cost compared to the current state-of-the-art models.
@@
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in vision monitoring applications of IoT devices, it is essential to transmit images of large areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for long-term monitoring, while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application levels, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy. However, to the best of our knowledge, this issue is largely unstudied for image transmission in existing deepJSCC literature.
@@
To address both efficiency and dynamic control over computational complexity, we propose feature importance-aware deepJSCC (FAJSCC). FAJSCC employs dimension-specialized computation strategies that enhance computational efficiency while achieving high transmission performance. The first strategy is axis-dimension specialized computation. A standard convolution over all channels, height, and width can be factorized into a depthwise convolution over the spatial dimensions (height and width) and a pointwise convolution over the channel dimension. This decomposition substantially reduces the number of operations: Instead of jointly processing all input and output channels with multi-channel kernels, the depthwise convolution handles each channel separately with each single-channel kernel, and the pointwise convolution mixes channels with 1×1 kernels. By exploiting this separation, FAJSCC achieves significantly lower computational cost while maintaining the feature processing capacity of standard convolution. To further enhance these operations, FAJSCC integrates spatial and channel attention mechanisms. The spatial attention emphasizes key features along the spatial axes, whereas the channel attention highlights important channels [35]. Accordingly, spatial attention is applied before depth-wise convolution and channel attention before point-wise convolution, maximizing feature processing for each axis. These lightweight attention operations exhibit linear computational complexity with respect to the input feature sizes. Our experiments demonstrate that even using only axis-dimension specialized computation can outperform the current state-of-the-art SwinJSCC [7] while using only half of the computational resources.
@@
Importantly, FAJSCC applies this deformable self-attention selectively based on feature importance. Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [39]. Less informative features can be processed sufficiently by the preceding axis-dimension specialized operations alone. Moreover, simple images require only minimal computation to achieve high task performance [40, 41]. Inspired by these findings, FAJSCC applies this deformable self-attention only to the most important features rather than to all features. This selective enhancement strategy greatly reduces computational cost compared with applying self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to the entire feature set. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity. Unlike previous lightweight deepJSCC methods that reduce or remove features [33, 30, 34], which degrade performance, FAJSCC enlarges and selectively refines important features, which increase performance while reducing computational burden.
@@
Moreover, FAJSCC decouples computational complexity from the model architecture, allowing first independent evaluation of the impact of encoder and decoder complexity on performance. It is important to note that typical previous deepJSCC models have fixed and symmetric computational complexity at encoder and decoder [5, 7, 12] or the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [30, 34]. As a result, such a discussion has not been possible in previous models. Recall that the encoder compresses images and adds redundancy to protect data against channel noise, while the decoder interprets the received noisy signal and (re)generates the transmitted images. A natural question arises: Which component requires the most computational complexity, and under what conditions? By independently varying the importance ratios for the encoder and decoder across different signal-to-noise ratios (SNRs), we draw two key conclusions: 1) the decoder requires higher computational complexity than the encoder, and 2) the decoder’s noisy-signal perception demands even greater complexity when the SNR is low. To the best of our knowledge, this is the first detailed analysis of computational complexity in deepJSCC.
@@
  • Adjustable Computation: We introduce the importance ratio to control the fraction of features processed by deformable self-attention, enabling FAJSCC to dynamically adjust computational complexity within a single model. Unlike previous methods that delete features to fit computational resources, FAJSCC preserves all information and allows independent adjustment of encoder and decoder complexity, making it possible to allocate computation according to available resources without compromising performance.
@@ II Problem Formulation
$\hat{y}_{i}=y_{i}+\epsilon_{i},~~\epsilon_{i}\sim\mathcal{CN}(0,\sigma^{2}),~i\in[1:k]$
where $\mathcal{CN}(0,\sigma^{2})$ is the circularly symmetric complex Gaussian distribution with mean 0 and variance $\sigma^{2}$. In the case of the fast Rayleigh fading channel with noise power $\sigma^{2}$, the relationship between the $i$-th symbols of $\mathbf{y}$ and $\hat{\mathbf{y}}$ is given by
$\hat{y}_{i}=h_{i}y_{i}+\epsilon_{i},~~h_{i}\sim\mathcal{CN}(0,1),~i\in[1:k]$
@@
${\textrm{SNR}}:=10\log_{10}\left(\frac{P}{\sigma^{2}}\right)=10\log_{10}\left(\frac{1}{\sigma^{2}}\right)~\text{dB}$
@@
After receiving $\hat{\mathbf{y}}\in\mathbb{C}^{k}$, the deepJSCC decoder $g_{\phi}$ with parameter $\phi$ maps the received signal $\hat{\mathbf{y}}$ to a reconstructed source image $\hat{\mathbf{x}}\in\mathbb{R}^{H\times W\times 3}$. Our goal is to develop a computationally efficient and adjustable deepJSCC model that minimizes the distortion between the original and reconstructed images, $\mathbf{x}$ and $\hat{\mathbf{x}}$, respectively. In this paper, the quality of reconstructed images is measured in two ways. First, peak signal-to-noise ratio (PSNR) measures the amount of pixel-wise distortion, defined in decibel (dB) as follows.
${\textrm{PSNR}}:=10\log_{10}\frac{{\textrm{MAX}}^{2}}{{\textrm{MSE}}}~\text{dB},$
@@
${\textrm{MSE}}:=\frac{1}{H\times W\times 3}\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2}.$
Second, the structural similarity index measure (SSIM) evaluates image quality based on human perception factors such as brightness, contrast, and structure [42]. Specifically, SSIM is defined by
${\textrm{SSIM}}(\mathbf{x},\hat{\mathbf{x}}):=l(\mathbf{x},\hat{\mathbf{x}})^{p}c(\mathbf{x},\hat{\mathbf{x}})^{q}s(\mathbf{x},\hat{\mathbf{x}})^{r},$
@@
To measure the efficiency of the deepJSCC model, we evaluate computational resource usage in terms of floating-point operations (FLOPs) and memory usage in bytes. FLOPs refer to the number of arithmetic operations involving floating-point numbers, such as additions and multiplications.
@@ III-B Selective Deformable Self-Attention
$\mathbf{Z}_{{\textrm{out}}}=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\intercal}}{\sqrt{C_{{\textrm{in}}}}}\right)\mathbf{V},$
@@
$(u,v)=(i,j)+\mathbf{z}_{{\textrm{offset}}}(i,j),\quad i_{0}=\lfloor u\rfloor,\;i_{1}=i_{0}+1,$
$j_{0}=\lfloor v\rfloor,\;j_{1}=j_{0}+1,\quad\alpha=u-i_{0},\;\;\beta=v-j_{0},$
$z_{{\textrm{out}}}^{c}(i,j)=\begin{bmatrix}1-\alpha&\alpha\end{bmatrix}\begin{bmatrix}z_{{\textrm{in}}}^{c}(i_{0},j_{0})&z_{{\textrm{in}}}^{c}(i_{0},j_{1})\\ z_{{\textrm{in}}}^{c}(i_{1},j_{0})&z_{{\textrm{in}}}^{c}(i_{1},j_{1})\end{bmatrix}\begin{bmatrix}1-\beta\\ \beta\end{bmatrix}$.
@@
Deformable self-attention effectively enhances the capability of standard self-attention. Building on this, we propose to applying deformable self-attention selectively to improve computational efficiency. Previous studies in neural source coding have shown that not all features carry the same amount of information: Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [39]. Moreover, it has been demonstrated that simple images do not require high computational cost to achieve high performance in super-resolution and image compression tasks [40, 47, 41]. Motivated by these observations, we apply deformable self-attention only to features in the most important windows, identified based on their contribution to image transmission performance. By restricting self-attention to these windows, the computational cost is significantly reduced. This selective enhancement strategy greatly reduces computational cost compared with applying deformable self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to features in all windows. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity.
@@
Figure 5 gives the detailed procedures of our proposed selective deformable self-attention. Firstly, based on the window importance $\mathbf{z}_{{\textrm{imp}}}\in[0,1]^{\frac{H_{{\textrm{in}}}}{\sqrt{n}}\times\frac{W_{{\textrm{in}}}}{\sqrt{n}}}$, windows are selected whether to be enhanced via deformable self-attention or not. The number of selected windows is determined by the importance ratio $\gamma$. When $\gamma=1$, features in all windows are processed via deformable self-attention, whereas $\gamma=0$ means none are processed. For $0<\gamma<1$, the features in the top $\gamma$ fraction windows, ranked by their measured importance, are selected and processed through deformable self-attention, which results in $\mathcal{O}\left(\frac{\gamma H_{{\textrm{in}}}W_{{\textrm{in}}}}{n}\left(4nC_{{\textrm{in}}}^{2}+2n^{2}C_{{\textrm{in}}}\right)\right)$ computational complexity. Secondly, the feature deformation process of Figure 4 is applied to the selected windows to adjust self-attention regions based on the input feature relations. Lastly, self-attention is applied to the selected adjusted areas and restore the original input feature shape with processed selected features and unselected features.
Unlike previous lightweight deepJSCC methods that reduce or remove features [33, 30, 34], which often degrade performance, FAJSCC selectively enlarges and refines important features, achieving higher performance while reducing computational cost. Moreover, by adjusting the importance ratio, FAJSCC decouples computational complexity from the model architecture, allowing first independent evaluation of the impact of encoder and decoder complexity on performance. The prior resource-adjustable methods [30, 34] cannot achieve this, as the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [30, 34]. Experiments varying encoder and decoder resources independently reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. Detailed results are shown in Section IV-B.

### v4

@@ Abstract
Recent advances in deep learning-based joint source–channel coding (deepJSCC) have substantially improved communication performance, but their growing computational cost hinders practical deployment. Moreover, certain applications require the ability to dynamically adapt computational complexity. To address these issues, we propose a Feature Importance-Aware deepJSCC (FAJSCC) model for image transmission that is both computationally efficient and adjustable. FAJSCC employs axis-dimension specialized computation, which performs efficient operations individually for each spatial and channel axis, significantly reducing computational cost while representing features effectively. It further incorporates selective deformable self-attention, which applies self-attention only to selected and adaptively adjusted features, leveraging the importance and relations of input features to efficiently capture complex feature correlations. Another key feature of FAJSCC is that the number of selected important areas can be controlled separately by the encoder and the decoder, depending on the available computational budget. It makes FAJSCC the first deepJSCC architecture to allow independent adjustment of encoder and decoder complexity within a single trained model. Experimental results show that FAJSCC achieves superior image transmission performance under various channel conditions while requiring less computational complexity than recent state-of-the-art models. Furthermore, experiments independently varying the encoder and decoder’s computational resources reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. The code is publicly available at github.com/hansung-choi/FAJSCCv2.
@@ I Introduction
Although conventional separation-based digital communication systems have undergone rapid improvements in both source coding (e.g., JPEG, JPEG2000, PNG, BPG, and AVIF) and channel coding (e.g., Turbo, LDPC, and polar codes) over several decades, further manual improvements to these rule-based methods have become increasingly difficult in satisfying recent communication requirements. In the other side, it is known that the joint source-channel coding (JSCC) design strictly outperforms the separation-based design in several scenarios, such as finite block length [1] and sending correlated data over multi-user settings [2]. Early analytic research on the JSCC was limited to simple data distributions, like Gaussian cases, due to the tractability, and hence could not deal with real data, such as images [3, 4].
@@
Beyond such compression-based strategies, another line of work focuses on designing lightweight models [29, 30, 31, 32, 33]. For example, reducing the intermediate deepJSCC feature channel size before passing through computationally intensive blocks [33] or performing computations only on a fixed subset of feature channels [29, 31] can lower the computational load. In addition, approaches that remove the least important features have been proposed [30, 34]. While these methods alleviate computational burden without requiring extra training, reducing or deleting features and restricting computations to limited channel regions degrade image transmission performance. Therefore, there is an urgent need for deepJSCC designs that can improve transmission performance while reducing computational cost compared to the current state-of-the-art models.
@@
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in vision monitoring applications of IoT devices, it is essential to transmit images of large areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for long-term monitoring, while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application levels, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy.
Although dynamic inference mechanisms such as routing, token pruning, and early-exit have been widely studied to adjust computational complexity in deep learning [35, 36, 37], they are not well suited to communication scenarios. Routing-based approaches require multiple models to be stored, leading to large memory overhead, which is impractical for memory-constrained IoT devices [38]. Token pruning-based designs reduce computational cost by discarding less important features [36], but this strategy is unsuitable for image transmission where all visual information must be preserved, and it also limits computational adaptability at the receiver side [30, 34]. Early-exit based designs such as DD-JSCC [39] adjust complexity by skipping intermediate layers, but the degree of layer skipping must be synchronized between the transmitter and receiver due to the dependency of feature distributions across layers. Consequently, the computational complexity is constrained by the side with stricter resource limitations and cannot be independently adjusted when the transmitter and receiver have heterogeneous computational capabilities. These limitations indicate that directly adopting existing dynamic inference frameworks is insufficient for deepJSCC systems, motivating the need for a communication-oriented design that enables independent computational resource adaptation within a single model.
@@
To simultaneously achieve computational efficiency and enable independent dynamic control of computational complexity at both the transmitter and receiver, we propose a feature importance-aware deepJSCC (FAJSCC) framework. FAJSCC employs dimension-specialized computation strategies that enhance computational efficiency while achieving high transmission performance. The first strategy is axis-dimension specialized computation. A standard convolution over all channels, height, and width can be factorized into a depthwise convolution over the spatial dimensions (height and width) and a pointwise convolution over the channel dimension. This decomposition substantially reduces the number of operations: Instead of jointly processing all input and output channels with multi-channel kernels, the depthwise convolution handles each channel separately with each single-channel kernel, and the pointwise convolution mixes channels with 1×1 kernels. By exploiting this separation, FAJSCC achieves significantly lower computational cost while maintaining the feature processing capacity of standard convolution. To further enhance these operations, FAJSCC integrates spatial and channel attention mechanisms. The spatial attention emphasizes key features along the spatial axes, whereas the channel attention highlights important channels [40]. Accordingly, spatial attention is applied before depth-wise convolution and channel attention before point-wise convolution, maximizing feature processing for each axis. These lightweight attention operations exhibit linear computational complexity with respect to the input feature sizes. Our experiments demonstrate that even using only axis-dimension specialized computation can outperform the current state-of-the-art SwinJSCC [7] while using only half of the computational resources.
@@
Importantly, FAJSCC applies this deformable self-attention selectively based on feature importance. Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [44]. Less informative features can be processed sufficiently by the preceding axis-dimension specialized operations alone. Moreover, simple images require only minimal computation to achieve high task performance [45, 46]. Inspired by these findings, FAJSCC applies this deformable self-attention only to the most important features rather than to all features. This selective enhancement strategy greatly reduces computational cost compared with applying self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature channel sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to the entire feature set. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity. Unlike previous lightweight deepJSCC methods that reduce or remove features [33, 30, 34], which degrade performance, FAJSCC enlarges and selectively refines important features, which increase performance while reducing computational burden. Moreover, due to our proposed selective feature enhancement framework, the transmitter and the receiver of FAJSCC can adjust their computational complexity independently, unlike previous resource-adjustable communication methods where resource adjustments are constrained by the other side [30, 34, 39].
@@
Moreover, FAJSCC decouples computational complexity from the model architecture, allowing first independent evaluation of the impact of encoder and decoder complexity on performance. It is important to note that typical previous deepJSCC models have fixed and symmetric computational complexity at encoder and decoder [5, 7, 12] or the transceiver’s computational complexity adjustments are constrained by the counterpart communication side [30, 34, 39]. As a result, such a discussion has not been possible in previous models. Recall that the encoder compresses images and adds redundancy to protect data against channel noise, while the decoder interprets the received noisy signal and (re)generates the transmitted images. A natural question arises: Which component requires the most computational complexity, and under what conditions? By independently varying the importance ratios for the encoder and decoder across different signal-to-noise ratios (SNRs), we draw two key conclusions: 1) the decoder requires higher computational complexity than the encoder, and 2) the decoder’s noisy-signal perception demands even greater complexity when the SNR is low. To the best of our knowledge, this is the first detailed analysis of computational complexity in deepJSCC.
@@
  • Adjustable Computation: We introduce the importance ratio to control the fraction of features processed by deformable self-attention, enabling FAJSCC to separately adjust the computational complexities of the encoder and the decoder. Unlike FAJSCC, previous resource-adjustable deep learning based communications are constrained by the counterpart communication side [30, 34, 39].
@@ II Problem Formulation
$\hat{y}_{i}=y_{i}+\epsilon_{i},~~\epsilon_{i}\sim\mathcal{CN}(0,\sigma^{2}),~i\in[1:k]$ (2)
where $\mathcal{CN}(0,\sigma^{2})$ is the circularly symmetric complex Gaussian distribution with mean 0 and variance $\sigma^{2}$.
In the case of the fast Rayleigh fading channel with noise power $\sigma^{2}$, the relationship between the $i$-th symbols of $\mathbf{y}$ and $\hat{\mathbf{y}}$ is given by
$\hat{y}_{i}=h_{i}y_{i}+\epsilon_{i},~~h_{i}\sim\mathcal{CN}(0,1),~i\in[1:k]$ (3)
@@
${\textrm{SNR}}:=10\log_{10}\left(\frac{P}{\sigma^{2}}\right)=10\log_{10}\left(\frac{1}{\sigma^{2}}\right)~\text{dB}$ (4)
@@
After receiving $\hat{\mathbf{y}}\in\mathbb{C}^{k}$, for the AWGN channel, the deepJSCC decoder $g_{\phi}$ with parameter $\phi$ maps the received signal $\hat{\mathbf{y}}$ to a reconstructed source image $\hat{\mathbf{x}}\in\mathbb{R}^{H\times W\times 3}$. In case of a fast Rayleigh fading channel, we assume perfect CSI only at the receiver side, i.e., Rayleigh fading coefficients are known to the receiver. We use the equalized signal $\bar{y}$ as decoder input with minimum mean square error equalization (MMSE), i.e., $\bar{\mathbf{y}}=\frac{\mathbf{h}^{*}}{\mathbf{h}^{*}\mathbf{h}+\sigma^{2}}\hat{\mathbf{y}}$. Here $\mathbf{h}^{*}$ is the complex conjugate of $\mathbf{h}$.
Our goal is to develop a computationally efficient and adjustable deepJSCC model that minimizes the distortion between the original and reconstructed images, $\mathbf{x}$ and $\hat{\mathbf{x}}$, respectively. This objective is particularly important because accurate image reconstruction is central to our target scenarios, such as wireless surveillance and monitoring applications. In this paper, the quality of reconstructed images is measured in two ways. Firstly, peak signal-to-noise ratio (PSNR) measures the amount of pixel-wise distortion, defined in decibel (dB) as follows.
${\textrm{PSNR}}:=10\log_{10}\frac{{\textrm{MAX}}^{2}}{{\textrm{MSE}}}~\text{dB},$ (5)
@@
${\textrm{MSE}}:=\frac{1}{H\times W\times 3}\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2}.$ (6)
Secondly, the structural similarity index measure (SSIM) evaluates image quality based on brightness, contrast, and structure [47]. Specifically, SSIM is defined by
${\textrm{SSIM}}(\mathbf{x},\hat{\mathbf{x}}):=l(\mathbf{x},\hat{\mathbf{x}})^{p}c(\mathbf{x},\hat{\mathbf{x}})^{q}s(\mathbf{x},\hat{\mathbf{x}})^{r},$ (7)
@@
To measure the efficiency of the deepJSCC model, we evaluate computational resource usage in terms of floating-point operations (FLOPs) and memory usage in bytes. FLOPs refer to the number of arithmetic operations involving floating-point numbers, such as additions and multiplications. We also measure latency (ms/image), peak memory (MB), and model storage size (MB). Here, latency is the end-to-end execution time from encoder to decoder, and peak memory is the maximum GPU memory allocation (MB) observed during a single forward pass of the model.
@@ III-B Selective Deformable Self-Attention
$\mathbf{Z}_{{\textrm{out}}}=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\intercal}}{\sqrt{C_{{\textrm{in}}}}}\right)\mathbf{V},$ (8)
@@
$(u,v)=(i,j)+\mathbf{z}_{{\textrm{offset}}}(i,j),\quad i_{0}=\lfloor u\rfloor,\;i_{1}=i_{0}+1,$ (9)
$j_{0}=\lfloor v\rfloor,\;j_{1}=j_{0}+1,\quad\alpha=u-i_{0},\;\;\beta=v-j_{0},$ (10)
$z_{{\textrm{out}}}^{c}(i,j)=\begin{bmatrix}1-\alpha&\alpha\end{bmatrix}\begin{bmatrix}z_{{\textrm{in}}}^{c}(i_{0},j_{0})&z_{{\textrm{in}}}^{c}(i_{0},j_{1})\\ z_{{\textrm{in}}}^{c}(i_{1},j_{0})&z_{{\textrm{in}}}^{c}(i_{1},j_{1})\end{bmatrix}\begin{bmatrix}1-\beta\\ \beta\end{bmatrix}$. (11)
@@
Deformable self-attention effectively enhances the capability of standard self-attention. Building on this, we propose to applying deformable self-attention selectively to improve computational efficiency. Previous studies in neural source coding have shown that not all features carry the same amount of information: Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [44]. Moreover, it has been demonstrated that simple images do not require high computational cost to achieve high performance in super-resolution and image compression tasks [45, 52, 46]. Motivated by these observations, we apply deformable self-attention only to features in the most important windows, identified based on their contribution to image transmission performance. By restricting self-attention to these windows, the computational cost is significantly reduced. This selective enhancement strategy greatly reduces computational cost compared with applying deformable self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature channel sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to features in all windows. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity.
@@
Figure 5 gives the detailed procedures of our proposed selective deformable self-attention. Firstly, based on the window importance $\mathbf{z}_{{\textrm{imp}}}\in[0,1]^{\frac{H_{{\textrm{in}}}}{\sqrt{n}}\times\frac{W_{{\textrm{in}}}}{\sqrt{n}}}$, windows are selected whether to be enhanced via deformable self-attention or not. The number of selected windows is determined by the importance ratio $\gamma$. When $\gamma=1$, features in all windows are processed via deformable self-attention, whereas $\gamma=0$ means none are processed. For $0<\gamma<1$, the features in the top $\gamma$ fraction windows, ranked by their measured importance, are selected and processed through deformable self-attention, which results in $\mathcal{O}\left(\frac{\gamma H_{{\textrm{in}}}W_{{\textrm{in}}}}{n}\left(4nC_{{\textrm{in}}}^{2}+2n^{2}C_{{\textrm{in}}}\right)\right)$ computational complexity. Secondly, the feature deformation process of Figure 4 is applied to the selected windows to adjust self-attention regions based on the input feature relations. Lastly, self-attention is applied to the selected adjusted areas, and the original input feature shape is restored with processed selected features and unselected features. During the restoration phase, the deformed features used in selective self-attention are mapped back to their original spatial positions by applying the inverse transformation of the feature deformation process.
Unlike previous lightweight deepJSCC methods that reduce or remove features [33, 30, 34], which often degrade performance, FAJSCC selectively enlarges and refines important features, achieving higher performance while reducing computational cost. Moreover, by adjusting the importance ratio, FAJSCC decouples computational complexity from the model architecture, allowing first independent evaluation of the impact of encoder and decoder complexity on performance. The prior resource-adjustable methods cannot achieve this, as the transceiver’s computational complexity adjustments are constrained by the counterpart communication side [30, 34, 39]. Experiments varying encoder and decoder resources independently reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. Detailed results are shown in Section IV-B.